OpenStudy (anonymous):

Sharmi and Reese each deposit money into a savings account. Sharmi deposits $200 less than Reese. After the same amount of time, Reese has earned $90 in interest at a rate of 3% per year, and Sharmi has earned $100 in interest at a rate of 5% per year. How much money did Reese and Sharmi initially deposit into their savings accounts?

OpenStudy (anonymous):

come on guys u can do it please help!

OpenStudy (anonymous):

simple interest or compound???

OpenStudy (anonymous):

idk what either of that is

OpenStudy (anonymous):

u should use the formula I= prt

OpenStudy (anonymous):

what is the formula is it \[\large{I= Prt}\] or, \[\large{A= P\times(1+r)^n}\]

OpenStudy (anonymous):

i was taught to use I = prt but whatever is easiest

OpenStudy (anonymous):

ok that is the formula for simple interest

OpenStudy (anonymous):

ok so shami and reese have deposited the amount for the same amount of time right

OpenStudy (anonymous):

uh yeah guess

OpenStudy (anonymous):

ok let us assume shami has deposited amount x so reese will deposit how much???

OpenStudy (anonymous):

x-200

OpenStudy (anonymous):

Sharmi deposits $200 less than Reese.

OpenStudy (anonymous):

correct your value

OpenStudy (anonymous):

it would be x+200

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Now in Shami's Case P = $ x r = 5% I = $ 100 Now plug in these values in this equation and solve for t \[\large{I= P\times{r}\times{t}}\]

OpenStudy (anonymous):

so 100= x(0.5) t ???

OpenStudy (anonymous):

no r = 5% = 5/100 = 0.05

OpenStudy (anonymous):

\[100= x(0.05) t\] ???

OpenStudy (anonymous):

yes now solve for t

OpenStudy (anonymous):

but the answer choices give two solutions

OpenStudy (anonymous):

ok wait first solve for t

OpenStudy (anonymous):

on solving for t you will get \[\large{t= {{100}\over{0.05x}}}= {{10000}\over{5x}}\]

OpenStudy (anonymous):

Similarly in reese's Case P = $ (x+200) r = 3% I = $ 90 Now plug in these values in this equation and solve for t \[\large{I= P\times{r}\times{t}}\]

OpenStudy (anonymous):

you will get \[\large{t= {{90}\over{0.03(x+200)}}}\] now equate these two equations

OpenStudy (anonymous):

ohhh okay so i solve for sharmi and then reese and thats the answer

OpenStudy (anonymous):

you will get \[\large{{10000}\over{5x}}= \large{{90}\over{0.03(x+200)}}\] now solve for x

OpenStudy (anonymous):

is it 600 ??

OpenStudy (anonymous):

cross multiply you will get 10000*0.03(x+200) = 5x*90 300(x+200) = 450x 300x + 60000 = 450x 450x -300x = 60000 150x = 60000 x = 60000/150 x = 400 so shami deposited $ 400 and reese deposited $400 + $200 = $600

OpenStudy (anonymous):

ok thank you so much!

OpenStudy (anonymous):

have u understood how to solve